# Relevance for Human Robot Collaboration

### v2

@@ V-E Relevance in Perception
By selectively processing the regions containing relevant elements, a notable maximum reduction of 30.09% in inference time and 26.53% in total time per frame of an object detector is achieved. Additionally, relevance improves two safety metrics by 11.25% and 13.50%, respectively [3].
@@ VI-B Demonstration Results
The demonstration results are shown in Fig. 4(b-h). In Fig. 4 (b), in the absence of a human in the scene, due to the lack of triggers, the perception module is continuously running, and no relevance determination is initialized. In Fig. 4(c-d), a human agent with recorded preferences for coffee enters the scene and grabs a coffee. The trigger is activated, and relevance is determined. Based on the recorded preferences, the robot actions are automatically generated to serve the human agent whole milk and a stir stick. In Fig. 4(e), a second human enters without indicative actions. The trigger is activated, but the relevance determination module lacks sufficient cues. The system continues looping through the perception module to collect more information. In Fig. 4(f-g), a conversation between two human agents about coffee activates the trigger to determine relevance. Without preference for the second human, the robot first generates actions related to the necessary elements (i.e., the coffee) and then inquires about the second human’s preferences for making the coffee. Based on the human response, the system generates a complete action sequence to serve the second human agent. In Fig. 4(h), two humans successfully get their preferred coffee with the robot’s assistance, demonstrating the effectiveness of relevance.
Figure 4: The illustration of (a) experimental setup and (b-h) demonstration results. With relevance, the robot fluently and successfully assists two human agents with cold-brew coffee drinking.
@@ VII Conclusion
This paper presents a novel concept and scene-understanding approach, termed ‘relevance,’ designed to identify pertinent elements within a scene to better support human objectives. To quantify relevance, we developed an innovative framework that continuously monitors the scene and selectively activates relevance assessment. We developed a probabilistic methodology with a novel scene representation to efficiently and accurately quantify relevance. Simulations demonstrate the framework’s effectiveness, achieving a precision of 0.99 and a recall of 0.94. to efficiently and accurately determine relevance. Relevance can significantly enhance HRC by improving task planning time by 79.56% compared to pure planning for a cereal task, reducing perception latency by up to 26.53% for object detection, enhancing safety by up to 13.50%, and decreasing the number of inquiries by 80.84%. The experimental results show a significant advancement toward creating a versatile robotic assistant capable of providing intelligent support to humans.

### v4

@@ V-E Relevance in Perception
Our previous work demonstrated that by selectively processing the regions containing relevant elements, a notable maximum reduction of 30.09% in inference time and 26.53% in total time per frame of an object detector is achieved. Additionally, relevance improves two safety metrics by 11.25% and 13.50%, respectively [7].
@@ VI-A Experimental Setup
The demonstration is implemented in Python using the threading package for multi-threading and asynchronous computation. The communication between threads is achieved using Event and Queue. At the beginning of the code, we initialized two threads, as shown in the perception module of Fig. 1, one for the visual information processing using the OpenAI API [35] and another one for picking up the microphone and processing the audio information using Assembly AI API [36]. In the visual information processing thread, we currently implement a VLM model for visual information processing, extracting the semantic information of all the objects on the table, classifying the objects into class and element representation, and detecting the human motion and objectives. For each iteration in the visual threads, the trigger criteria are checked. We implement the trigger criterion, the changes in human numbers in the scene, which is sufficient for our current demonstration. Once a trigger criterion is met, a new thread for relevance determination is initialized, and relevance is determined based on the methodologies in Section IV. The probabilities required for the computation are derived from an LLM model in the AI toolkit. We currently leverage the OpenAI API, and the model we used is GPT-4o [35]. Within the same thread, the action is generated based on the relevance identified to best assist humans.
@@ VI-B Demonstration Results
The demonstration results are shown in Fig. 4(b-h). In Fig. 4 (b), in the absence of a human in the scene, due to the lack of triggers, the perception module is continuously running, and no relevance determination is initialized. In Fig. 4(c-d), a human agent with recorded preferences for coffee enters the scene and grabs a coffee. The trigger is activated, and the human’s objective and relevance are determined based on the methodologies we described in Section IV. Based on the recorded preferences, the robot actions are automatically generated to serve the human agent whole milk and a stir stick. In Fig. 4(e), a second human enters without indicative actions. The trigger is activated, but the relevance determination module lacks sufficient cues. The system continues looping through the perception module to collect more information. In Fig. 4(f-g), a conversation between two human agents about coffee provides more cues to determine relevance. Without preference for the second human, the robot first generates actions related to the necessary elements (i.e., the coffee) and then inquires about the second human’s preferences for making the coffee. Based on the human response, the system generates a complete action sequence to serve the second human agent. In Fig. 4(h), two humans successfully get their preferred coffee with the robot’s assistance, demonstrating the effectiveness of relevance. Without the unique components in relevance, such as the evaluation of cue sufficiency, a flexible probabilistic framework integrating all the factors, a novel hierarchical scene representation, etc., this type of HRC assistance in a general setup will be challenging. The demonstration video can be found here: https://www.youtube.com/watch?v=XVUxWbwsYhs
Figure 4: The illustration of (a) experimental setup and (b-h) demonstration results. With relevance, the robot successfully and seamlessly assists two human agents with cold-brew coffee drinking.
@@ VII Conclusion
This paper presents a novel concept, dimensionality reduction approach, and HRC framework termed ‘relevance’. Relevance is a dimensionality reduction process that incorporates a continuously operating perception module, evaluates cue sufficiency within the scene, and applies a flexible formulation and computation framework considering diversified factors. The perception module runs continuously and leverages multiple modalities of sensors and processing to provide cues and information. The event-based framework has the capability to determine if the relevance determination should be triggered. A probabilistic methodology to evaluate relevance based on a novel hierarchical scene representation can efficiently and accurately quantify relevance, considering various human and environmental factors. Simulations demonstrate the framework’s effectiveness, achieving a precision of 0.99, a recall of 0.94, an F1 score of 0.96, and an object ratio of 0.94. Relevance can significantly enhance HRC by accurately and effectively improving task planning time by 79.56% compared to pure planning for a cereal task, reducing perception latency by up to 26.53% for object detection, enhancing safety by up to 13.50%, and decreasing the number of inquiries by 80.84%. The experimental demonstration of serving two people coffee drinking shows that relevance enables the robots to assist humans in a proactive, seamless, and structured manner. This is a significant advancement toward creating a versatile robotic assistant capable of providing intelligent support to humans.